OpenStudy (anonymous):

Select the correct graph of -x - 2y = 6. Answer graph of line crossing the y axis at 3 and through point parentheses 1 comma 1 close parentheses graph of line crossing the y axis at negative 3 and through point parentheses 2 comma negative 4 close parentheses graph of line crossing the y axis at 3 and through point parentheses 2 comma 2 close parentheses graph of line crossing the y axis at negative 3 and through point parentheses 1 comma negative 5 close parentheses

OpenStudy (anonymous):

-x -2y = 6 -2y = 6 + x 2y = 6 - x y = -(1/2)x + 3 y intercept is at 3 and crosses through (2, 2)

OpenStudy (anonymous):

Thanks !